OpenStudy (anonymous):

Help for a MEDAL and FAN!!!!! A magnetic hockey puck (0.40 kg) hits another hockey puck (0.40 kg) at rest and both travel together to the right at 8.0 m/s. How fast was the first hockey puck initially moving?

OpenStudy (caozeyuan):

well, you have to notice that the total momentum of the system does not change. so \[m _{1}v _{1}=m _{2}v _{2}\]

OpenStudy (caozeyuan):

you have m1=0.4kg, m2=o.4+0.4=o.8kg, and v2=8m*s^-1

OpenStudy (caozeyuan):

so v1=0.8*8/0.4=16 meter per second

OpenStudy (anonymous):

Ooh, Okay. Thank you!

OpenStudy (anonymous):

A 6000 kg train car traveling at 2.5 m/s collides with another 6000 kg train car, now they are both moving but with a speed of 1.25 m/s. This is an example of

OpenStudy (caozeyuan):

do you mean filling int he blank after " of “ ?